OpenStudy (anonymous):

Find the rate of change for this graph. How did you find the rate of change?

OpenStudy (anonymous):

http://prntscr.com/23xxev

OpenStudy (anonymous):

dy/dx change in y = dy = (25-0) change in x = dx = (3-0) dy/dx = 25/3

OpenStudy (anonymous):

Thank you so much ^-^